### v3

# Do Automatic Factuality Metrics Measure Factuality? A Critical Evaluation

Abstract
Modern LLMs can now produce highly readable abstractive summaries, to the point where traditional automated metrics for evaluating summary quality, such as ROUGE, have become saturated. However, LLMs still sometimes introduce unwanted content into summaries, i.e., information inconsistent with or unsupported by their source. Measuring the occurrence of these often subtle “hallucinations” automatically has proved to be challenging. This in turn has motivated development of a variety of metrics intended to measure the factual consistency of generated summaries against their source. But are these approaches measuring what they purport to do? In this work, we stress-test automatic factuality metrics. Specifically, we investigate whether and to what degree superficial attributes of summary texts suffice to predict “factuality”, finding that a (supervised) model using only such shallow features is reasonably competitive with SOTA factuality scoring methods. We then evaluate how factuality metrics respond to factual corrections in inconsistent summaries and find that only a few show meaningful improvements. In contrast, some metrics are more sensitive to benign, non-factual edits. Motivated by these insights, we show that one can “game” (most) automatic factuality metrics, i.e., reliably inflate “factuality” scores by appending innocuous sentences to generated summaries. Taken together, our results raise questions about the degree to which we should rely on existing automated factuality metrics and what exactly we want “factuality metrics” to measure.
Do Automatic Factuality Metrics Measure Factuality?
A Critical Evaluation
Sanjana Ramprasad Northeastern University ramprasad.sa@northeastern.edu                        Byron C. Wallace Northeastern University b.wallace@northeastern.edu
Figure 0.1: Many methods have been proposed to automatically evaluate the faithfulness (or factual consistency) of summaries with respect to inputs. In this work we critically evaluate such approaches, e.g., by measuring their sensitivity to various manipulations, as shown here.
1 Introduction
LLMs are generally strong abstractive summarizers Goyal et al. (2022); Zhang et al. (2024). Still, they are not infallible. Even the largest, most capable models sometimes introduce subtle “hallucinations” (or “confabulations”) into summaries that are unsupported by or in contradiction to the corresponding input document (Zhang et al., 2024; Tang et al., 2024b; Ramprasad et al., 2024b). Such behaviour is especially problematic in domains such as medicine or law, where inaccurate information could translate into meaningfully negative consequences for individuals.
However, manually evaluating model outputs’ faithfulness with respect to their source is expensive, time-consuming, and impractical to scale. This has motivated the development of automated methods that score generated summaries for faithfulness with respect to reference documents. Such metrics have been operationalized using a range of underlying techniques, including entailment (Laban et al., 2022; Scirè et al., 2024), QA models (Scialom et al., 2021; Fabbri et al., 2021b), specialized models explicitly trained to score source-summary pairs (Zhong et al., 2022; Zha et al., 2023; Tang et al., 2024a), and more recently, prompt-based methods that rely on LLM calls (Luo et al., 2023; Wang et al., 2023a).
Most of these metrics have been evaluated against human benchmark assessments (binary labels or Likert scores) of factual consistency (Maynez et al., 2020; Fabbri et al., 2021a; Laban et al., 2022; Honovich et al., 2022; Wang et al., 2022; Gao et al., 2023; Tang et al., 2024a). These assessments have established that automated factuality metrics correlate with human evaluations to varying degrees. But are such approaches actually attuned to subtle question of factual consistency between inputs and outputs, or are they merely relying on shallow heuristic signals—such as lexical overlap, entity repetition, or the presence of specific phrases—to make their judgments? For instance, Kamoi et al. (2023b) assessed the reliability of QA-based metrics and found that while they can predict summary-level factuality, they fail to identify correct error spans, and are outperformed by simple exact match baselines.
If metrics do rely on shallow heuristic patterns, this raises serious concerns about the reliability of these metrics, making them vulnerable to “gaming”, where summaries or claims can be crafted to exploit heuristic shortcuts and artificially inflate scores without genuinely improving factual accuracy.
In this work, we stress test a diverse set of SOTA factual consistency metrics on context-summary pairs that yield continuous factual consistency scores. Our analyses offer the following empirical results as our main contributions.
Superficial attributes of text predict “factuality” nearly as well as complex SOTA metrics when supervised. We define a small set of straightforward heuristic predictors (e.g. lexical overlap, conciseness, word novelty etc.) By themselves, such features do not directly imply factual consistency (or a lack thereof). But we find that fine-tuning a simple MLP defined over these features typically yields performance competitive with (often complex) SOTA factual consistency metrics. This suggests that it is at least possible that existing metrics rely (merely) on superficial cues which correlate with consistency, rather than making a nuanced assessment of accuracy.
Factuality metrics are somewhat responsive to factual corrections, but also often sensitive to irrelevant (benign) modifications to summaries. We use a previously released dataset (Krishna et al., 2024) comprising a small set of summaries labelled as being inconsistent with the corresponding input, paired with minimally edited versions that fix these inconsistencies. We use this set to assess whether factuality metrics are responsive to these edits, i.e., whether the scores they assign to the faithful versions of summaries are higher than the original inconsistent versions. We find that most metrics are at least somewhat responsive in this way. However, we also consider the complement of this experiment: We modify summaries in ways that should not affect factual consistency (e.g., paraphrasing texts) and measure changes in factuality scores assigned. Metrics would ideally be invariant to such benign perturbations, but we find that they are, on average, sensitive to them. In fact, some metrics exhibit greater responsiveness to benign changes in inaccurate summaries than to actual factual corrections, raising concerns about their reliability in distinguishing genuine improvements from inconsequential edits.
(Some) factual consistency metrics are gameable. The preceding findings suggest that factuality metrics rely at least partially on superficial cues that do not directly reflect factual consistency. In turn, this suggests that we should be able to “game” them by inserting cues to systematically inflate scores assigned to model outputs. Indeed, we find that superfluous innocuous phrases used standalone or with summaries induce higher factuality scores. For many fine-tuned metrics, these artificial boosts from “gamed” summaries exceed the gains achieved from summaries generated by more factual models, exposing a critical vulnerability.
While previous efforts have meta-evaluated factuality metrics (Gabriel et al., 2021; Chen et al., 2021a), these approaches primarily focus on inducing errors to assess metrics’ sensitivity to specific error types or their general detection capabilities. In contrast, our work examines the degree to which factuality metrics respond to factual changes compared to unrelated textual attributes or artifacts, shedding light on their susceptibility to spurious correlations and non-factual influences.
2 Experimental Setup
2.1 Benchmark Datasets
Automatic factuality metrics for summarization have been evaluated using various benchmarks, predominantly based on news sources (Hermann et al., 2015; Narayan et al., 2018). The AggreFact dataset (Tang et al., 2022) consolidates several benchmarks on fine-tuned model-generated summaries from such sources (Maynez et al., 2020; Wang et al., 2020; Pagnoni et al., 2021; Fabbri et al., 2021a; Honovich et al., 2022; Laban et al., 2022). Similarly, datasets like FacEval (Wang et al., 2022) and ReferenceMatters (Gao et al., 2023) benchmark dialogue summarization of fine-tuned models on sources from SAMSum (Gliwa et al., 2019a) and DialogSum (Chen et al., 2021b).
However, these datasets primarily focus on summaries generated by fine-tuned models that predate modern LLMs, which have demonstrated superior zero-shot summarization capabilities (Goyal et al., 2022). As a result, the performance of factuality metrics on these benchmarks may not accurately reflect their effectiveness on outputs from SOTA LLMs, which are prone to different types of errors (Tang et al., 2022). To bridge this gap, Tang et al. (2024a) introduced the LLM-AggreFact dataset, which includes factuality labels for summaries generated by recent LLMs across a variety of domains, including Wikipedia, interviews, news, dialogues, science, and healthcare. Similarly, Krishna et al. (2024) released the GenAudit dataset with factuality annotations for LLM summaries in news, Reddit, and clinical settings, while Ramprasad et al. (2024a) focus on LLM generated dialogue summaries in the LLM-dialogue dataset.
For our analysis we use all of the above benchmarks to capture a wide range of error types. For fine-tuned model summaries, we use AggreFact for news and FacEval for dialogues. For LLM-generated summaries, we rely on LLM-AggreFact, GenAudit, and LLM-dialogue across varied domains. We note that each benchmark consolidates multiple datasets and to ensure clean separation of distributions, we avoid overlapping datasets between our test and development splits. We achieve this by randomly selecting datasets from each benchmark for the development set (we include only the dev split), while the remaining datasets are reserved for testing (using their respective test split). Given its smaller size, the GenAudit dataset is included exclusively in the test set. A detailed breakdown of the dev and test splits is provided in Appendix A.1.
2.2 Automatic Factuality Metrics for Summarization
We group SOTA factuality metrics into four broad methodological categories. This includes metrics based on: Question Answering (QA), Natural Language Inference (NLI), fine-tuned (specialized) models and LLM prompting methods; see Table 1.
For QA-based metrics we use QuestEval (Scialom et al., 2021). As an NLI-based metric we use SummaC-Conv (Laban et al., 2022). We use three specialized models for evaluation: UniEval, AlignScore and MiniCheck. UniEval Zhong et al. (2022) reframes NLG evaluation as a Boolean QA task and uses T5 (Raffel et al., 2020) to score different dimensions. AlignScore (Zha et al., 2023) evaluates summaries by combining an alignment function—a RoBERTa model (Liu, 2019) fine-tuned on diverse tasks—with a splitting and aggregation strategy. MiniCheck(Tang et al., 2024a) uses a Flan-T5 model (Chung et al., 2022) fine-tuned on a synthetic dataset created by the authors.
We also use GPT-4 to score the factual consistency of summaries based on a direct assessment (DA) prompt template from Wang et al. (2023a).
Metric Category
QuestEval (Scialom et al., 2021) QA
SummaC-Conv (Laban et al., 2022) NLI
UniEval (Zhong et al., 2022) specialized model
AlignScore (Zha et al., 2023) specialized model
MiniCheck (Tang et al., 2024a) specialized model
ChatGPT-DA (Wang et al., 2023a) Prompt / LLM
Table 1: Metrics categorized by approach and analyzed
2.3 Heuristic Indicators of “Factuality”
The methods for automatically scoring summary factuality reviewed above often rely on complex models, such as QA or NLI-based approaches, or specialized models. Could simpler heuristics provide comparable signal? Here we explore a set of simple features for predicting “faithfulness” to evaluate if they are sufficient for this task.
One of the simplest features we use is lexical overlap between a summary and its source. We measure this using ROUGE-2 F1 which matches word pairs or bigrams between the summary and the source.
Figure 2.1: We train a simple MLP (bottom bar) over only superficial indicators of consistency (e.g., word overlap). This realizes performance comparable to SOTA consistency metrics for in most cases. This model is supervised, but we train and test on disjoint datasets. Our results raise questions about how much inference is necessary to assess consistency on these benchmark corpora, given what superficial predictors reveal.
We also include entity overlap, i.e., the proportion of entities in the summary that are present in the corresponding source text.
Another predictor we use is semantic similarity. Specifically, embed summary and source sentences via BERT (Devlin, 2018), then score each summary sentence based on which source sentence it is most similar to (using cosine similarity). We average all summary sentence scores to derive a final score.
We include word and sentence novelty ratio features, which measure the proportion of words and sentences in the summary that do not appear in the source, respectively.
Finally, we measure text conciseness by calculating a conciseness ratio, or the ratio of the length (number of words) of the source document to the length of the summary. This measures the extent to which a summary has condensed the original text.
3 Are Superficial Features Sufficient to Infer “Factuality”?
The degree to which a summary is faithful to the text it summarizes is a subtle question that demands understanding the content in both texts to determine if they are consistent. We would therefore expect that metrics capable of measuring factual consistency in general would need to capitalize on information beyond the superficial features of source and summary texts just discussed.
To assess if this is in fact the case we first evaluate the extent to which a model using superficial features alone can be trained to infer factuality. We train an MLP to predict (human) factuality labels as a function of shallow input features. We compare the results achieved—specifically, AUC scores—using this method to a suite of more sophisticated approaches that have been previously proposed.
The shallow model has been explicitly supervised. To replicate the scenario of metrics trained on different datasets, we ensure that our training set is entirely separate from the benchmark datasets in the test set; this makes the approach akin to the supervised, “specialized” models for measuring factual consistency considered.
We emphasize that the purpose of this exercise is not to propose a practical “new” approach to score summaries for factuality. Rather, our point here is to determine the extent to which existing metrics improve performance beyond what can be achieved using simple, superficial attributes of text.
Figure 2.1 summarizes our findings. Perhaps surprisingly, this shallow model is generally competitive with some previously proposed (complex) models. On document and query-focused summarization (QFS), the shallow model performs comparably to established QA/NLI metrics and the custom metric UniEval. In fact, it outperforms these metrics in the context of dialogue summarization evaluation. However, the shallow model trails Alignscore, MiniCheck, and the ChatGPT-DA metrics, suggesting that these may be more nuanced. Overall, our results establish that one can achieve performance that is at least generally competitive with SOTA factuality consistency metrics using only information captured by shallow features. [^1] This does not imply that existing metrics are overly or exclusively relying on such cues, only that they might.
4 What do automatic factuality scores measure?
We next attempt to characterize what factuality metrics measure. First, we examine the correlations between shallow features and the scores produced by these metrics (4.1). Then (4.2), we evaluate the sensitivity of the metrics to controlled changes in summary faithfulness, measuring how effectively they respond to corrections of factual inaccuracies. This allows us to assess their ability to detect and reflect improvements in factual consistency.
4.1 Predicting Automatic Factuality Measures with Shallow Features
Figure 4.1: Spearman correlations between actual and predicted metric scores from an MLP model trained to replicate each metric using shallow features as inputs. Results show moderate-to-strong correlations, indicating potential reliance on superficial text attributes.
We aim to evaluate the extent to which metrics rely on shallow features. To this end we first assess whether we can replicate the scores assigned to summaries by SOTA factuality consistency metrics using shallow predictors only.
To explore this, we again train an MLP model that takes these shallow features as inputs. However, instead of predicting factual inconsistency labels directly (as in the preceding section), we train the model to predict the corresponding scores assigned by the SOTA factuality metrics. In other words, we ask whether such metrics could, in principle, be a function solely of heuristic indicators. Figure 4.1 reports Spearman rank correlations between the predicted scores from the MLP and the metric scores on the test set.
These results show a moderate-to-strong correlation between the predicted scores of the MLP model, trained on shallow features, and the actual metric scores. Note, UniEval and ChatGPT exhibit the weakest correlation among the metrics. These correlations alone do not provide definitive evidence that the metrics necessarily over-rely on shallow indicators; however, they do suggest that such reliance may be a possibility. To try and establish a causal link (i.e., whether such metrics are sensitive to heuristic indicators rather than underlying changes in factual inconsistencies) we next perform a series of controlled manipulations to summaries and measure their responsiveness to these.
4.2 Measuring Metric Sensitivities to Controlled Manipulations
Here we examine the sensitivity of automatic factuality metrics to factual corrections and benign modifications. We aim to clarify whether metrics are responsive to changes in factual consistency, and/or if they are unduly influenced by superficial adjustments with no bearing on factual consistency.
Figure 4.2: Score differences for each metric between the original summary and summaries edited for factual accuracy by humans (shown in green) and other benign edits (shown in blue).
For this analysis, we focus on a subset of summaries manually labelled as inconsistent by human annotators. For each of these, we have a paired, minimally edited version that has been corrected for consistency. These inconsistent/corrected summary pairs were released with the Genaudit dataset (Krishna et al., 2024). Using these, we can measure metric sensitivity to factual consistency (the actual quality of interest).
We are also interested in score variation that owes to superfluous factors. To this end, we prompt GPT-4 to create versions of summaries modified to vary in targeted ways that should be independent of factual consistency. For this we use several prompts requesting different transformations to summaries: Adding least relevant sentences from the source to extend summaries (added source text); Reducing vocabulary diversity (less diverse); Introducing negation words in a meaning-preserving way (negated); Paraphrasing; Shuffling sentence order (shuffled), and; Simplifying text (simplified) or replacing random words in the summary with a synonym (synonym replacement). We provide prompts for summary edits in Appendix C.1.
Ideally, automatic factuality metrics should show a positive score change for corrected summaries, reflecting improved factual consistency. Conversely, generated summary rewrites—which do not alter factual content—should exhibit minimal score changes. Any significant score differences in these rewrites likely indicate that the metric is sensitive to artifacts incidental to factual consistency.
Results
We present our results in Figure 4.2. Specialized models MiniCheck and AlignScore, and prompt-based ChatGPT-DA show the strongest positive score shifts following factual corrections, as one would hope for: Corrected, more factual summaries are rated higher than inconsistent ones. SummaC-Conv and UniEval also show positive, albeit smaller, differences. QuestEval shows no change (on average) for corrected summaries.
Most metrics end up showing significant score increases when a random source sentence is inserted into an inconsistent summary—often surpassing or matching the gains from factual corrections. [^2] Such a sentence will itself be “consistent”, but should probably not shift the overall consistency score upwards given that this does not correct any existing inconsistencies, as it strictly adds more content. QuestEval and SummaC-Conv exhibit the largest increases in this dimension, while AlignScore and UniEval show comparable improvements to corrected scores. In contrast, MiniCheck and ChatGPT-DA exhibit positive shifts but smaller than those observed for corrected summaries.
Another clear pattern emerges across all metrics, especially with MiniCheck, UniEval, and ChatGPT-DA show a decrease in scores when meaning-preserving negation terms are introduced. These negations—added through simple rephrasing without changing the original meaning (e.g., changing “The author mistakenly believed their ACT test was on a certain day” to “The author did not realize their ACT test was not on the day they thought”)—suggest that these models are particularly sensitive to negation.
SummaC and UniEval are more sensitive to benign modifications than to corrections. In contrast, AlignScore and MiniCheck exhibit greater sensitivity across a range of benign modifications. While ChatGPT is sensitive to negated sentences, it tends to be more robust to other types of modifications.
Taken together, these results demonstrate the variability of factual consistency metrics with respect to spurious and meaningful edits to outputs. And given that at least some of the metrics appear to be responsive to inconsequential edits, our results imply that we might be able to game metrics, i.e., systematically inflate scores without attendant changes in factual consistency. [^3] One could argue that such manipulations—adversarial ones especially—result in “out of distribution” inputs, and so we should have no expectation of how models will perform. But usually factual consistency metrics are touted (implicitly) as measuring consistency between arbitrary document and summary candidate pairs.
5 Can we game factuality metrics?
Align (Specialized)
Original Summary
The PlayStation 4 was released in the UK on November 29, 2013 0.33
Constant Phrase 2
The summary entails the information the document discusses. 0.93
Manipulated Summary
The PlayStation 4 was released in the UK on November 29, 2013. The summary entails the information the document discusses. 0.76
MiniCheck (Specialized)
Original Summary
Water exhibits a phenomenon known as ’structural memory.’ 0.005
Constant Phrase 1
The document discusses 0.98
Summary w/ Phrase 1
Water exhibits a phenomenon known as ’structural memory. The document discusses. 0.49
Table 2: Qualitative (cherry-picked) samples of original and manipulated summaries with corresponding metric scores for AlignScore and MiniCheck. For comprehensiveness, we report quantitative aggregated results in Figure 5.1, and we provide more examples in Appendix 7.
Figure 5.1: Pairwise score differences across summaries manipulated using four distinct strategies: Adding constant strings (top phrase and assertion phrase) and appending them to summaries (summ + top, and summ + assertion. The results reveal that NLI and bespoke model metrics are particularly vulnerable to gaming, with significant score inflation observed under these manipulations.
Given the preceding observations—which suggest that superficial cues may influence automated factuality metrics, we turn to a practical question that would exploit this behavior: Are these metrics gameable? In other words, can we systematically manipulate summaries to induce higher factuality scores? If one can reliably do this, it may raise concerns about the use of these metrics for tasks like leaderboard rankings. Independent of practical considerations, establishing the gameability of factuality metrics would provide additional evidence that they may not be measuring what we think.
As a “gaming” strategy, we introduce innocuous phrases containing no factual content, testing their effect on metric scores as standalone inputs and as additions to existing summaries. These phrases, when appended, should not alter the factual accuracy of the summaries they accompany. Our goal here is to assess whether such neutral modifications can nonetheless inflate factuality scores.
Strings for summary manipulation We attempt to identify a set of strings that artificially inflate metric scores. To do so, we first identify the top 20th percentile of summaries ranked by each factuality metric, reasoning that these high-scoring summaries may contain patterns that raise scores, independent of context. To this end, we compute TF-IDF scores for all bigrams in these summaries, identifying those disproportionately present in summaries deemed “factual” by each metric.
We select from these the top 100 bigrams, aggregating results across metrics. This set includes, e.g., “the document” and “document discusses”. We adopt the constant phrase “the document discusses” (top phrase) into all documents and observe whether metrics are responsive to this. We consider one additional phrase: such as “The summary entails information in the document.” (assertion phrase) which explicitly asserts factual consistency. The wording of this phrase varies slightly across metrics to align with their specific methods for evaluating factual consistency. The complete list of these phrases for each metric is provided in the Appendix D.1. Finally, we append both phrases to the corresponding summaries summ (+top) and summ (+assertion).
Our goal here is to determine whether the appended text unduly influences metrics, potentially resulting in inflated scores that do not accurately reflect the factual quality of the original content.
We report the average pairwise difference in metric scores between the gamed versions of summaries and their original versions when evaluated for consistency with the source. Examples of these manipulated summaries and their corresponding scores are provided in Table 2.
Results and Discussion
Metric sensitivity Figure 5.1 reports the effects of our gaming strategies on metric scores. Notably, the constant phrases boosts scores by >>>0.2 points (absolute) for NLI-based SummaC-Conv, as well as specialized models UniEval, AlignScore, and MiniCheck. This is surprising, as these phrases are not valid summaries and do not contain any factual content (note that top phrase is an incomplete sentence). Adding constant phrases as suffixes to summaries increased scores by 0.1–0.15 points; this is comparable to the gains realized following factual corrections, as reported in Section 4.2. In fact, SummaC-Conv shows no score increase for corrected summaries, suggesting an under-sensitivity to actual changes in factuality.
To further contextualize these numbers, we analyze metric score differences between summaries generated by larger models (e.g., GPT-4, Gemini) and smaller models (e.g., Llama-7B, Mistral-7B, Falcon-7B, BART). Larger models generally produce more consistent summaries (Tang et al., 2022; Goyal et al., 2022), providing a baseline for assessing gaming strategies. A natural question, then, is whether “gaming” achieves score improvements comparable to those that have resulted from advances in model size and complexity.
We analyze a subset of 115 document-summary pairs from GenAudit (Krishna et al., 2024) and Aggrefact-old (Tang et al., 2022), where model information is available. Our findings reveal that gaming with constant phrases leads to metric score improvements 3x to 8x greater than those achieved through genuine improvements in generated summaries, particularly for NLI and specialized metrics. For instance, gaming increases SummaC scores by nearly 0.2 points, while model improvements have minimal impact. Similarly, adding suffixes to summaries achieves gains comparable to those won with better summarization models. (Full results in Appendix D.1.)
QuestEval is mostly unaffected by these manipulations. ChatGPT-DA shows a negative trend: Scores drop when summaries are “gamed”. Further analysis suggests that the QA-based metric generates questions based on innocuous phrases and entity repetitions that introduce noise, leading to unchanged or lower scores. ChatGPT-DA’s performance decline likely results from its strict adherence to the prompts. A qualitative analysis, where the model was prompted to explain its scoring, revealed that in several instances, it explicitly identified the gaming phrase as misleading. However, we observe that ChatGPT-DA’s scores can be inflated by replacing summaries with a disclaimer acknowledging potential errors in interpreting the source (see Appendix D.2).
Overall, these results suggest that adding fixed phrases to summaries can boost metric scores—for most finetuned model metrics—at levels comparable to, or even exceeding, presumably genuine improvements due to advances in summarization models themselves.
Filler phrases and metric performance
The innocuous, non-factual phrases used in our experiments above constitute “filler” phrases, which introduce no substantive content. Since most factuality metrics primarily assess factual precision, we hypothesize that they may struggle to differentiate between informative content and irrelevant filler text leading to a weakness that could be exploited. Our experiments above provide some evidence of this and we explore it further in this section. First, in addition to “gaming prefixes” we consider two other types of constant strings that are neutral phrases: They are independent of our analysis of top-scoring bigrams and do not allude to the consistency of summaries with respect to the document. We measure the pairwise difference in metric scores when these phrases are appended as a point of reference.
We consider two neutral phrase types: a “baseline” phrase, consisting of two generic phrases that don’t mention the summary or source, and a “qualifier” phrase, suggesting the summary may have multiple interpretations (see Table 6 for phrases). This contrasts with our “assertion” phrase, which directly affirms the summary’s faithfulness. These filler phrases should ideally not affect consistency scores, as they don’t alter factual content. We also compare how metrics respond to these phrases versus “gaming” phrases to determine if only targeted fillers increase scores.
As shown in Appendix D.2, all metrics assign lower scores to summaries with a baseline phrase, indicating that not all irrelevant sentences inflate scores. However, qualifier phrases have similar effects to gaming prefixes on MiniCheck and UniEval, suggesting that neutral phrases referencing the summary may boost scores for these metrics. For SummaC and AlignScore, only specific gaming prefixes increase scores, while other phrases reduce them. Interestingly, ChatGPT, which usually shows a negative trend, exhibits inflated scores when the summary is replaced with a neutral qualifier phrase acknowledging potential errors in interpreting the source.
Although human evaluators may accept such neutral phrases as consistent, they likely assess models more comprehensively, identifying deeper inconsistencies. In contrast, metrics are not sophisticated enough to capture these nuances and can be influenced by irrelevant information, inflating scores independent of factual consistency.
6 Discussion
Distilling salient information from large volumes of unstructured text—i.e., summarization—is a compelling practical use case for LLMs. As they have grown increasingly competent at this task, it has become concomitantly difficult to evaluate their performance, which requires assessing often subtle but potentially critical factual inconsistencies between inputs and generated summaries. Safe deployment of, and continued progress on, such approaches to summarization requires automatic metrics to assess factual consistency automatically. Many such approaches have been proposed in the literature, but there has been little effort to stress-test them, which is the goal of the present work.
We found first that a simple MLP fine-tuned over superficial features and summaries can realize performance competitive with SOTA factual consistency metrics. We then showed that while such metrics are generally responsive to manual changes affecting factual consistency, they are also often sensitive to spurious transformations which should not affect scores. Motivated by these findings, we then showed that factual consistency metrics are often gameable, i.e., in many cases we can add fixed phrases as suffixes into output summaries or use them standalone to reliably inflate “factuality” scores beyond gains observed via improvements in model size and complexity.
Limitation and Ethics
This work has several important limitations that should be taken into account when interpreting our results.
First, in Section 3, we demonstrated that a model trained solely on heuristic indicators of factual consistency can achieve performance on par with state-of-the-art (SOTA) approaches, though this was a supervised model specifically for consistency detection. Despite our training data being sourced from entirely different datasets and domains than those used for evaluation, it is important to highlight that the dataset was manually annotated. However, we cannot yet conclude how a shallow model would perform if trained on heuristic-based predictors of synthetically generated summaries—this would be most similar to the data that other (specialized) metrics have been trained with. We emphasize, however, that our goal with these experiments was not to introduce a new metric, but rather to investigate how much current metrics improve upon basic textual attributes.
Another limitation here is that one could quibble with just how “shallow” these predictors are. Our view is that even the more involved features we use for this—e.g., semantic similarity based on BERT embeddings and entity overlap—should not be taken as reliable measures of factual consistency, but empirically they do correlate with manual annotations of factual consistency (see Appendix B). We can only speculate regarding the underlying causal structure based on these results.
Our subsequent interventions in Section 4.2 established that in some cases metrics are sensitive to extraneous transformations (e.g., re-phrasings). However this was not observed uniformly, and we performed such transformations using LLMs, which may have introduced legitimate factual inconsistencies (though we believe this to be rare). A limitation to our “gaming” experiment is that it involved a manual process; however, this does not affect the main result, which is that there exist strings that can be appended to outputs in such a way that consistently boosts scores.
A further limitation to this (and all manipulation experiments) is that they produce outputs which may be viewed as “out of distribution”; but generally the purpose of summary factuality metrics is to be able to apply them to arbitrary model outputs and corresponding references, so it is not entirely clear what “in distribution” means under this assumption.
Despite these limitations, we believe that—taken together—our results suggest that one should be careful about their interpretation of automatic factual consistency metrics. And we hope these results inform research on summarization moving forward.
Appendix A Experimental Setup
A.1 Benchmark Datasets
We list the datasets used in the benchmarks and specify which split (train/test) they are included in Table 3.
Benchmark Train Datasets Test Datasets
LLM-Aggrefact (Tang et al., 2024a) TofuEval-MeetB (Tang et al., 2024b),
ExpertQA (Malaviya et al., 2023)
Lfqa (Chen et al., 2023)
RAGTruth (Wu et al., 2023)
FactCheck-GPT (Wang et al., 2023b)
Wice (Kamoi et al., 2023a)
TofuEval-MediaS (Tang et al., 2024b)
ClaimVerify (Liu et al., 2023)
Reveal (Jacovi et al., 2024)
Aggrefact (Tang et al., 2022) FRANK (Pagnoni et al., 2021)
Polytope (Huang et al., 2020)
FactCC (Kryscinski et al., 2020)
Goyal21 (Goyal and Durrett, 2021)
Wang20 (Wang et al., 2020)
Cao22 (Cao et al., 2022)
XSumFaith (Maynez et al., 2020)
CLIFF (Cao and Wang, 2021)
SummEval (Fabbri et al., 2021a)
FacEval (Wang et al., 2022) SAMSUM (Gliwa et al., 2019b) -
LLM-Dialogue (Ramprasad et al., 2024a) SAMSUM DialogSum (Chen et al., 2021b)
Genaudit (Krishna et al., 2024) - REDDIT,
ACIBENCH (Yim et al., 2023),
XSUM (Narayan et al., 2018)
Table 3:
Appendix B Metric correlation with shallow features
Figure B.1: Spearman rank correlations between each automatic evaluation metric and individual shallow features. Some of these correlations are moderate, suggesting that in aggregate superficial features may carry much of the signal being measured by automated measures of factual consistency.
Rewrite type Prompt
Shuffled Rewrite the following text by changing the order of sentences without altering the original meaning of the text.
Note: You must not omit any information from the original text or alter its meaning.
Text: <summary>
Rewritten Text:
Added Source Text Edit the following summary by adding a sentence from the source. Ensure the source sentence added is the most irrelevant to the summary.
Source: <source>.
Text: <summary>
Edited Text:
Less
Diverse
Rewrite the following summary by decreasing the variety of vocabulary used.
Note: You must not omit any information from the original text or alter its meaning
Text: <summary>
Rewritten Text:
Negated Rewrite the following text by introducing negation in a way that preserves the original meaning of the text.
Note: You must not omit any information from the original text or alter its meaning
Text: <summary>
Rewritten Text:
Simplified Rewrite the following text by simplifying any complex sentences.
Note: You must not omit any information from the original text or alter its meaning
Text: <summary>
Shortened Text:
Shortened Rewrite the text concisely.
Note: You must not omit any information from the original text or alter its meaning
Text: <summary>
Rewritten Text:
Paraphrased Paraphrase the text.
Note: You must not omit any information from the original text or alter its meaning
Text: <summary>
Paraphrased Text:
Synonym
Replacement
Rewrite the following text by replacing some common words with a synonym.
Note: You must not omit any information from the original text or alter its meaning
Text: <summary>
Rewritten Text:
Table 4: List of summary manipulations along with the GPT-4 prompts used to rewrite summaries and obtain targetted manipulations.
Figure B.1 reports the Spearman rank correlation between each automatic evaluation metric and all shallow features. We observe a moderate correlation between many (metric, heuristic) pairs. For instance, most metrics show moderately negative correlations with word novelty. This suggests that metrics generally penalize summaries that introduce new vocabulary not present in the input text. [^4] Our paraphrasing experiments in Section 4.2 provide some further evidence for this; some (though not all) metrics are sensitive to paraphrasings of summaries which should not affect underlying factual consistencies. While this heuristic may help identify unsupported information, it also indicates a bias towards summaries that closely mirror the source language.
QA metrics (QuestEval, QAFactEval) and the prompt-based ChatGPT-DA exhibit the highest correlations with bigram overlap, showing moderate positive correlations (>>>0.4) while displaying weak correlations (0.2-0.4) with ROUGE and entity overlap. This bias may be beneficial in domains like scientific or legal documents, where exact phrasing is important. However, over-reliance on exact matches may be misleading in certain contexts.
The NLI-based metric SummaC-Conv, bespoke model metric and ChatGPT-DA show stronger stronger correlation with “semantic similarity” (under BERT) than bigram overlap (ROUGE), suggesting they may be better suited to evaluating abstractive summaries.
NLI metric SummaC exhibits the strongest positive correlation with conciseness while Bespoke model and prompt-based metrics show weaker correlations. This could indicate a potential preference for shorter summaries. In contrast, QA and prompt-based metrics show lower correlations with QAFactEval showing negligible correlations. This discrepancy highlights concerns about the fairness of comparing metrics across summaries of varying lengths, as some metrics may inherently favour brevity, which is unlikely to be causally linked to consistency.
Appendix C Measuring Metric Sensitivities to Controlled Manipulations
C.1 Summary Edits
Table 4 lists the prompts used to generate summary variants, each designed to vary in specific ways without affecting factual consistency.
Appendix D Can we game factuality metrics?
D.1 Metric-Specific Gaming Phrases
We provide the constant phrases (cost phrase 2) used per metric in Table 5
Metric Constant phrase 2
QuestEval The summary is consistent with the information the document discusses.
SummaC-Conv The summary entails the information the document discusses.
UniEval The claim is consistent with the information the document discusses.
AlignScore The summary entails the information the document discusses.
MiniCheck TThe claim entails the information the document discusses.
ChatGPT-DA The summary is consistent with the information the document discusses.
Table 5: Metrics and the corresponding constant phrase 2 used to game them.
Filler Type Filler phrase
Baseline In any case, understanding complex topics requires a multifaceted approach.
Qualifier This summary reflects one possible understanding, though interpretations may differ.
Table 6: Additional filler phrases used to artificially inflate scores
D.2 Metric Score Shifts: Gaming Strategies vs. Model Improvements
To contextualize results better, we show score improvements when summaries are gamed along with score improvements brought about by summaries from more complex models in Figure D.1
Figure D.1: Score improvements from gaming vs from using “better” models. Gaming strategies lead to score improvements comparable to or greater than boosts from model improvements.
Figure D.2: Score improvements from gaming with two other types of prefixes: “baseline” and “qualifier” compared to our “top” and “assertion” gaming prefixes. Baseline prefixes are not meaningful, while the qualifier prefix implies the summary’s accuracy can vary based on interpretation.
D.3 Gamed summary examples
Gamed summary examples are provided in Table 7
SummaC-Conv (NLI)
Original Summary
147 people, including 142 students, are in critical condition. 0.26
Constant Phrase 2
The summary entails the information the document discusses 0.99
Summary w/ Phrase 2
147 people, including 142 students, are in critical condition. The summary entails the information the document discusses 0.85
UniEval (Bespoke)
Original Summary
Byte Pair Encoding offers benefits in terms of confidentiality 0.48
Constant Phrase 2
The claim is consistent with the information the document discusses 0.99
Summary w/ Phrase 2
Byte Pair Encoding offers benefits in terms of confidentiality. The claim is consistent with the information the document discusses. 0.74
Align (Bespoke)
Original Summary
The PlayStation 4 was released in the UK on November 29, 2013 0.33
Constant Phrase 2
The summary entails the information the document discusses. 0.93
Manipulated Summary
The PlayStation 4 was released in the UK on November 29, 2013. The summary entails the information the document discusses. 0.76
MiniCheck (Bespoke)
Original Summary
Water exhibits a phenomenon known as ’structural memory.’ 0.005
Constant Phrase 1
The document discusses 0.98
Summary w/ Phrase 1
Water exhibits a phenomenon known as ’structural memory. The document discusses. 0.49
Table 7: Selected examples of original and manipulated summaries with corresponding metric scores. Only metrics identified as gameable are shown.
Appendix E Is the GPT-Based Prompt Approach Truly Grounded in the Source Document for Inconsistency Detection?
A key question when using GPT for document-based consistency checks is whether the model performs a true "closed-book" evaluation using the source document as context. Since GPT has access to a vast amount of prior knowledge, it raises the concern of whether it is verifying the summary based solely on this knowledge, without actually considering the source article for context. To assess how the presence or absence of contextual grounding impacts the scoring, we compare summary scores using two approaches: “without context”, where GPT is asked to rate summaries on a scale of 1 to 100 for factual accuracy without any reference to the source article, and “with context”, where GPT evaluates consistency relative to a source article using the Direct Assessment (DA) prompt.
Figure E.1: GPT scores of summaries with and without context.
Figure E.1 presents the distribution of GPT-generated scores under the two conditions Our analysis suggests that GPT’s reliance on internal knowledge versus provided context varies significantly depending on the task domain. The most significant differences emerge in the QFS domain, where the score distribution indicates that a larger proportion of summaries in the moderate range (0.4 - 0.6) are rated lower when GPT is provided with contextual information from the source document, possibly because inconsistencies went undetected without a grounding document. In contrast, the dialogue domain presents an opposite trend. When GPT is given access to the source context, more summaries receive higher scores, possibly because the conversational nature of the domain makes consistency harder to evaluate without context.
However, the overall similarity between the distributions across both conditions raises concerns. The fact that GPT appears to evaluate factual accuracy similarly, regardless of whether it has access to the original document, implies a potential overreliance on its internal knowledge, which could lead to misleading consistency scores. This could be particularly problematic in cases where the document contains information that conflicts with GPT’s internal knowledge. In such scenarios, GPT might incorrectly rate summaries as factually accurate because it defaults to what it “knows” rather than what the document states. This could undermine the reliability of GPT as an evaluator of document-grounded consistency and raise broader concerns about its suitability for tasks requiring strict adherence to external sources.
The implications are particularly significant for tasks involving dynamically updated or specialized knowledge, where prompting GPT could fail to recognize novel or contradictory information if it leans too heavily on pre-existing internal knowledge.

### v4

@@
 Abstract
-Modern LLMs can now produce highly readable abstractive summaries, to the point where traditional automated metrics for evaluating summary quality, such as ROUGE, have become saturated.
-However, LLMs still sometimes introduce unwanted content into summaries, i.e., information inconsistent with or unsupported by their source.
-Measuring the occurrence of these often subtle “hallucinations” automatically has proved to be challenging.
-This in turn has motivated development of a variety of metrics intended to measure the factual consistency of generated summaries against their source.
-But are these approaches measuring what they purport to do?
-In this work, we stress-test automatic factuality metrics.
-Specifically, we investigate whether and to what degree superficial attributes of summary texts suffice to predict “factuality”, finding that a (supervised) model using only such shallow features is reasonably competitive with SOTA factuality scoring methods.
-We then evaluate how factuality metrics respond to factual corrections in inconsistent summaries and find that only a few show meaningful improvements.
-In contrast, some metrics are more sensitive to benign, non-factual edits.
-Motivated by these insights, we show that one can “game” (most) automatic factuality metrics, i.e., reliably inflate “factuality” scores by appending innocuous sentences to generated summaries.
-Taken together, our results raise questions about the degree to which we should rely on existing automated factuality metrics and what exactly we want “factuality metrics” to measure.
-Do Automatic Factuality Metrics Measure Factuality?
-A Critical Evaluation
-Sanjana Ramprasad Northeastern University ramprasad.sa@northeastern.edu                        Byron C. Wallace Northeastern University b.wallace@northeastern.edu
-Figure 0.1: Many methods have been proposed to automatically evaluate the faithfulness (or factual consistency) of summaries with respect to inputs.
+Modern LLMs can now produce highly readable abstractive summaries, to the point that traditional automated metrics for evaluating summary quality, such as ROUGE, have saturated.
+However, LLMs still sometimes introduce inaccuracies into summaries, i.e., information inconsistent with or unsupported by the corresponding source.
+Measuring the occurrence of these often subtle factual inconsistencies automatically has proved challenging.
+This in turn has motivated development of metrics intended to measure the factual consistency of generated summaries against sources.
+But are these approaches measuring what they purport to?
+Or are they mostly exploiting artifacts?
+In this work, we stress test a range of automatic factuality metrics, including specialized models and LLM-based prompting methods, to probe what they actually capture.
+Using a shallow classifier to separate “easy” examples for factual evaluation where surface features suffice from “hard” cases requiring deeper reasoning, we find that all metrics show substantial performance drops on the latter.
+Furthermore, some metrics are more sensitive to benign, fact-preserving edits than to factual corrections.
+Building on this observation, we demonstrate that most automatic factuality metrics can be gamed, i.e., their scores can be artificially inflated by appending innocuous, content-free sentences to summaries.
+Among the metrics tested, the LLM-based ChatGPT-DA approach is the most robust and reliable.
+However, this comes with a notable caveat: Prompting LLMs to assess factuality may overly rely on their parametric knowledge rather than the provided reference when making judgments.
+Taken together, our findings call into question the reliability of current factuality metrics and prompt a broader reflection on what these metrics are truly measuring.
+We conclude with concrete recommendations for improving both benchmark design and metric robustness, particularly in light of their vulnerability to superficial manipulations.
+1 Introduction
+LLMs are strong abstractive summarizers (goyal2022news; zhang2024benchmarking), but they are not infallible.
+Even the largest, most capable models sometimes introduce subtle “hallucinations” (or “confabulations”) into summaries that are unsupported by or in contradiction to the corresponding input document (zhang2024benchmarking; tang-etal-2024-tofueval; ramprasad-etal-2024-evaluating).
+Such behavior is especially problematic in domains such as medicine or law, where inaccurate information could translate into meaningfully negative consequences for individuals.
+However, manually evaluating model outputs’ factual consistency with respect to references is expensive, time-consuming, and impractical to scale.
+This has motivated development of automated methods that score generated summaries for consistency with respect to reference documents.
+Such metrics have been operationalized using a range of techniques, including entailment (SummaC; laban2022summac), QA models (QuestEval; scialom2021questeval), specialized models explicitly trained to score source-summary pairs (UniEval; zhong2022towards, Alignscore; zha2023alignscore, MiniCheck; tang2024minicheck), and more recently, LLM-based methods that rely on prompting LLMs (ChatGPT-DA; wang2023chatgpt; luo2023chatgpt).
+Figure 1: Many methods have been proposed to automatically evaluate the factual consistency of summaries with respect to inputs.
 In this work we critically evaluate such approaches, e.g., by measuring their sensitivity to various manipulations, as shown here.
-1 Introduction
-LLMs are generally strong abstractive summarizers Goyal et al.
-(2022); Zhang et al.
-(2024).
-Still, they are not infallible.
-Even the largest, most capable models sometimes introduce subtle “hallucinations” (or “confabulations”) into summaries that are unsupported by or in contradiction to the corresponding input document (Zhang et al., 2024; Tang et al., 2024b; Ramprasad et al., 2024b).
-Such behaviour is especially problematic in domains such as medicine or law, where inaccurate information could translate into meaningfully negative consequences for individuals.
-However, manually evaluating model outputs’ faithfulness with respect to their source is expensive, time-consuming, and impractical to scale.
-This has motivated the development of automated methods that score generated summaries for faithfulness with respect to reference documents.
-Such metrics have been operationalized using a range of underlying techniques, including entailment (Laban et al., 2022; Scirè et al., 2024), QA models (Scialom et al., 2021; Fabbri et al., 2021b), specialized models explicitly trained to score source-summary pairs (Zhong et al., 2022; Zha et al., 2023; Tang et al., 2024a), and more recently, prompt-based methods that rely on LLM calls (Luo et al., 2023; Wang et al., 2023a).
-Most of these metrics have been evaluated against human benchmark assessments (binary labels or Likert scores) of factual consistency (Maynez et al., 2020; Fabbri et al., 2021a; Laban et al., 2022; Honovich et al., 2022; Wang et al., 2022; Gao et al., 2023; Tang et al., 2024a).
+Most of these metrics have been evaluated against human benchmark assessments (binary labels or Likert scores) of factual consistency (maynez-etal-2020-faithfulness; fabbri2021summeval; laban2022summac; honovich2022true; wang2022analyzing; gao2023reference; tang2024minicheck).
 These assessments have established that automated factuality metrics correlate with human evaluations to varying degrees.
-But are such approaches actually attuned to subtle question of factual consistency between inputs and outputs, or are they merely relying on shallow heuristic signals—such as lexical overlap, entity repetition, or the presence of specific phrases—to make their judgments?
-For instance, Kamoi et al.
-(2023b) assessed the reliability of QA-based metrics and found that while they can predict summary-level factuality, they fail to identify correct error spans, and are outperformed by simple exact match baselines.
-If metrics do rely on shallow heuristic patterns, this raises serious concerns about the reliability of these metrics, making them vulnerable to “gaming”, where summaries or claims can be crafted to exploit heuristic shortcuts and artificially inflate scores without genuinely improving factual accuracy.
-In this work, we stress test a diverse set of SOTA factual consistency metrics on context-summary pairs that yield continuous factual consistency scores.
+But are such approaches actually attuned to the subtle question of factual consistency between inputs and outputs, or are they merely relying on shallow heuristic signals to make their judgments?
+If metrics do rely on shallow heuristic patterns, this raises questions about their reliability.
+It also makes them vulnerable to “gaming”: Summaries or claims can be crafted to exploit heuristic shortcuts and artificially inflate scores without genuinely improving factual accuracy.
+Alternatively, LLM-based approaches may overly rely on their internal (parametric) knowledge to assess summaries, rather than assessing output factuality with respect to the source.
+In this work, we stress test a diverse set of SOTA factual consistency metrics on context-claim pairs that yield continuous factual consistency scores.
 Our analyses offer the following empirical results as our main contributions.
-Superficial attributes of text predict “factuality” nearly as well as complex SOTA metrics when supervised.
-We define a small set of straightforward heuristic predictors (e.g.
-lexical overlap, conciseness, word novelty etc.) By themselves, such features do not directly imply factual consistency (or a lack thereof).
-But we find that fine-tuning a simple MLP defined over these features typically yields performance competitive with (often complex) SOTA factual consistency metrics.
-This suggests that it is at least possible that existing metrics rely (merely) on superficial cues which correlate with consistency, rather than making a nuanced assessment of accuracy.
+Metrics struggle when factuality requires reasoning beyond shallow cues.
+In Section 3 we train a shallow MLP on surface-level heuristics and use its prediction confidence to categorize summaries as easy, moderate, or hard to assess for factuality.
+We find that all automated metrics exhibit substantial performance drops from easy to hard examples, suggesting they may be relying on superficial cues rather than nuanced reasoning.
 Factuality metrics are somewhat responsive to factual corrections, but also often sensitive to irrelevant (benign) modifications to summaries.
-We use a previously released dataset (Krishna et al., 2024) comprising a small set of summaries labelled as being inconsistent with the corresponding input, paired with minimally edited versions that fix these inconsistencies.
-We use this set to assess whether factuality metrics are responsive to these edits, i.e., whether the scores they assign to the faithful versions of summaries are higher than the original inconsistent versions.
-We find that most metrics are at least somewhat responsive in this way.
-However, we also consider the complement of this experiment: We modify summaries in ways that should not affect factual consistency (e.g., paraphrasing texts) and measure changes in factuality scores assigned.
-Metrics would ideally be invariant to such benign perturbations, but we find that they are, on average, sensitive to them.
-In fact, some metrics exhibit greater responsiveness to benign changes in inaccurate summaries than to actual factual corrections, raising concerns about their reliability in distinguishing genuine improvements from inconsequential edits.
+In Section 4.1, we evaluate whether factuality metrics distinguish genuine factual corrections from inconsequential edits using a dataset of annotated summary pairs (krishna2024genaudit) which comprise an inconsistent summary and a minimally edited, faithful revision.
+Ideally, metrics should assign higher scores to the corrected versions while remaining stable under benign, fact-preserving edits (e.g., paraphrasing).
+We find that while both specialized and prompt-based metrics respond to factual corrections, many—especially NLI-based SummaC and some specialized models (UniEval and Alignscore)—are overly sensitive to benign edits.
+In fact, some metrics exhibit greater score shifts from superficial changes than from actual factual improvements.
+By contrast, ChatGPT-DA consistently ranks faithful revisions higher while remaining robust to benign perturbations, suggesting that it may be a more reliable choice.
 (Some) factual consistency metrics are gameable.
-The preceding findings suggest that factuality metrics rely at least partially on superficial cues that do not directly reflect factual consistency.
-In turn, this suggests that we should be able to “game” them by inserting cues to systematically inflate scores assigned to model outputs.
-Indeed, we find that superfluous innocuous phrases used standalone or with summaries induce higher factuality scores.
-For many fine-tuned metrics, these artificial boosts from “gamed” summaries exceed the gains achieved from summaries generated by more factual models, exposing a critical vulnerability.
-While previous efforts have meta-evaluated factuality metrics (Gabriel et al., 2021; Chen et al., 2021a), these approaches primarily focus on inducing errors to assess metrics’ sensitivity to specific error types or their general detection capabilities.
-In contrast, our work examines the degree to which factuality metrics respond to factual changes compared to unrelated textual attributes or artifacts, shedding light on their susceptibility to spurious correlations and non-factual influences.
+If factuality metrics rely at least partially on superficial cues, this suggests that we should be able to “game” them by inserting such cues to inflate scores assigned to model outputs.
+And indeed in Section 5 we find that inserting superfluous, innocuous phrases—either alone or appended to claims——can significantly inflate factuality scores.
+For NLI-based and specialized metrics, these artificial boosts often exceed the score gains achieved by genuinely more factual models (see Figure 6).
+ChatGPT-DA, however, shows minimal sensitivity to such manipulations, again suggesting that prompting LLMs for factuality assessment may offer more robust and reliable evaluations.
+However, this comes with a caveat.
+GPT sometimes over relies on internal knowledge when evaluating summaries—even when a reference source is provided.
+ChatGPT-DA may default to its internal (parametric) knowledge rather than relying on the provided source when assessing factuality consistency (see Section 4.2).
+This tendency appears especially pronounced when the reference document contradicts the model’s parametric knowledge, indicating that its reliability may be compromised in cases involving myths, rare or updated facts, or information that conflicts with what the model “believes”.
+This provides a caveat to our other findings that suggest LLM based assessments may be more reliable than bespoke models: LLM-based evaluations may reflect what the (evaluator) model “knows” rather than what the source says, limiting reliability in grounded factuality assessments.
+Practical recommendations.
+Our analyses reveal that while specialized metrics perform well overall, they are vulnerable to benign edits and adversarial manipulation, which may limit their reliability.
+ChatGPT-DA offers a more robust alternative, but may emphasize consistency with its parametric knowledge rather than input sources, raising concerns about its grounding behavior.
+We therefore recommend caution when applying such metrics in domains involving myths, misinformation, or uncommon facts.
+Finally, we highlight the need for benchmarks that capture hallucination severity and for methods that incorporate saliency-aware supervision to improve metric reliability (Section 7)
 2 Experimental Setup
 2.1 Benchmark Datasets
-Automatic factuality metrics for summarization have been evaluated using various benchmarks, predominantly based on news sources (Hermann et al., 2015; Narayan et al., 2018).
-The AggreFact dataset (Tang et al., 2022) consolidates several benchmarks on fine-tuned model-generated summaries from such sources (Maynez et al., 2020; Wang et al., 2020; Pagnoni et al., 2021; Fabbri et al., 2021a; Honovich et al., 2022; Laban et al., 2022).
-Similarly, datasets like FacEval (Wang et al., 2022) and ReferenceMatters (Gao et al., 2023) benchmark dialogue summarization of fine-tuned models on sources from SAMSum (Gliwa et al., 2019a) and DialogSum (Chen et al., 2021b).
-However, these datasets primarily focus on summaries generated by fine-tuned models that predate modern LLMs, which have demonstrated superior zero-shot summarization capabilities (Goyal et al., 2022).
-As a result, the performance of factuality metrics on these benchmarks may not accurately reflect their effectiveness on outputs from SOTA LLMs, which are prone to different types of errors (Tang et al., 2022).
-To bridge this gap, Tang et al.
-(2024a) introduced the LLM-AggreFact dataset, which includes factuality labels for summaries generated by recent LLMs across a variety of domains, including Wikipedia, interviews, news, dialogues, science, and healthcare.
-Similarly, Krishna et al.
-(2024) released the GenAudit dataset with factuality annotations for LLM summaries in news, Reddit, and clinical settings, while Ramprasad et al.
-(2024a) focus on LLM generated dialogue summaries in the LLM-dialogue dataset.
+Automatic factuality metrics for fact verification have been evaluated using various benchmarks, predominantly based on news sources (hermann2015teaching; narayan-etal-2018-dont).
+The AggreFact dataset (tang2022understanding) consolidates several benchmarks on fine-tuned model-generated summaries from such sources (maynez-etal-2020-faithfulness; wang2020asking; pagnoni2021understanding; fabbri2021summeval; honovich2022true; laban2022summac).
+Similarly, datasets like FacEval (wang2022analyzing) and ReferenceMatters (gao2023reference) benchmark dialogue summarization of fine-tuned models on sources from SAMSum (gliwa-etal-2019-samsum) and DialogSum (chen-etal-2021-dialogsum).
+These datasets primarily focus on summaries from fine-tuned models predating modern LLMs, which excel in zero-shot summarization (goyal2022news).
+Consequently, factuality metrics evaluated on these benchmarks may not generalize to SOTA LLM outputs, which exhibit different error patterns (tang2022understanding).
+To address this, tang2024minicheck introduced LLM-AggreFact, a fact-checking dataset from recent LLMs across diverse domains and dataset types.
+Similarly, krishna2024genaudit released the GenAudit dataset with factuality annotations for LLM summaries in news, Reddit, and clinical settings, while ramprasad-etal-2024-analyzing focus on LLM generated dialogue summaries in the LLM-dialogue dataset.
 For our analysis we use all of the above benchmarks to capture a wide range of error types.
 For fine-tuned model summaries, we use AggreFact for news and FacEval for dialogues.
-For LLM-generated summaries, we rely on LLM-AggreFact, GenAudit, and LLM-dialogue across varied domains.
+For LLM-generated summaries, we rely on LLM-AggreFact, GenAudit, and LLM-dialogue.
 We note that each benchmark consolidates multiple datasets and to ensure clean separation of distributions, we avoid overlapping datasets between our test and development splits.
-We achieve this by randomly selecting datasets from each benchmark for the development set (we include only the dev split), while the remaining datasets are reserved for testing (using their respective test split).
-Given its smaller size, the GenAudit dataset is included exclusively in the test set.
-A detailed breakdown of the dev and test splits is provided in Appendix A.1.
+Since AggreFact primarily includes older fine-tuned models, we incorporate its dev set into our development data and reserve newer benchmarks for evaluation.
+Specifically, our dev set includes summaries from the AggreFact dev split, as well as XSUM and CNNDM examples from Genaudit, ensuring no overlap with test data.
+All remaining datasets are evaluated using their respective test splits.
+We provide a detailed breakdown of our dev and test splits in Appendix 8.
 2.2 Automatic Factuality Metrics for Summarization
@@
 This includes metrics based on: Question Answering (QA), Natural Language Inference (NLI), fine-tuned (specialized) models and LLM prompting methods; see Table 1.
-For QA-based metrics we use QuestEval (Scialom et al., 2021).
-As an NLI-based metric we use SummaC-Conv (Laban et al., 2022).
+For QA-based metrics we use QuestEval (scialom2021questeval).
+As an NLI-based metric we use SummaC-Conv (laban2022summac).
 We use three specialized models for evaluation: UniEval, AlignScore and MiniCheck.
-UniEval Zhong et al.
-(2022) reframes NLG evaluation as a Boolean QA task and uses T5 (Raffel et al., 2020) to score different dimensions.
-AlignScore (Zha et al., 2023) evaluates summaries by combining an alignment function—a RoBERTa model (Liu, 2019) fine-tuned on diverse tasks—with a splitting and aggregation strategy.
-MiniCheck(Tang et al., 2024a) uses a Flan-T5 model (Chung et al., 2022) fine-tuned on a synthetic dataset created by the authors.
-We also use GPT-4 to score the factual consistency of summaries based on a direct assessment (DA) prompt template from Wang et al. (2023a).
+UniEval zhong2022towards reframes NLG evaluation as a Boolean QA task and uses T5 (raffel2020exploring) to score different dimensions.
+AlignScore (zha2023alignscore) evaluates summaries by combining an alignment function—a RoBERTa model (liu2019roberta) fine-tuned on diverse tasks—with a splitting and aggregation strategy.
+MiniCheck(tang2024minicheck) uses a Flan-T5 model (chung2022scaling) fine-tuned on a synthetic dataset created by the authors.
+We also use GPT-4o-mini to score the factual consistency of summaries based on a direct assessment (DA) prompt template from wang2023chatgpt.
 Metric Category
-QuestEval (Scialom et al., 2021) QA
-SummaC-Conv (Laban et al., 2022) NLI
-UniEval (Zhong et al., 2022) specialized model
-AlignScore (Zha et al., 2023) specialized model
-MiniCheck (Tang et al., 2024a) specialized model
-ChatGPT-DA (Wang et al., 2023a) Prompt / LLM
+QuestEval (scialom2021questeval) QA
+SummaC-Conv (laban2022summac) NLI
+UniEval (zhong2022towards) specialized model
+AlignScore (zha2023alignscore) specialized model
+MiniCheck (tang2024minicheck) specialized model
+ChatGPT-DA (wang2023chatgpt) Prompt / LLM
 Table 1: Metrics categorized by approach and analyzed
-2.3 Heuristic Indicators of “Factuality”
-The methods for automatically scoring summary factuality reviewed above often rely on complex models, such as QA or NLI-based approaches, or specialized models.
-Could simpler heuristics provide comparable signal?
-Here we explore a set of simple features for predicting “faithfulness” to evaluate if they are sufficient for this task.
-One of the simplest features we use is lexical overlap between a summary and its source. We measure this using ROUGE-2 F1 which matches word pairs or bigrams between the summary and the source.
-Figure 2.1: We train a simple MLP (bottom bar) over only superficial indicators of consistency (e.g., word overlap).
-This realizes performance comparable to SOTA consistency metrics for in most cases.
-This model is supervised, but we train and test on disjoint datasets.
-Our results raise questions about how much inference is necessary to assess consistency on these benchmark corpora, given what superficial predictors reveal.
-We also include entity overlap, i.e., the proportion of entities in the summary that are present in the corresponding source text.
-Another predictor we use is semantic similarity.
-Specifically, embed summary and source sentences via BERT (Devlin, 2018), then score each summary sentence based on which source sentence it is most similar to (using cosine similarity).
-We average all summary sentence scores to derive a final score.
-We include word and sentence novelty ratio features, which measure the proportion of words and sentences in the summary that do not appear in the source, respectively.
-Finally, we measure text conciseness by calculating a conciseness ratio, or the ratio of the length (number of words) of the source document to the length of the summary.
-This measures the extent to which a summary has condensed the original text.
-3 Are Superficial Features Sufficient to Infer “Factuality”?
-The degree to which a summary is faithful to the text it summarizes is a subtle question that demands understanding the content in both texts to determine if they are consistent.
-We would therefore expect that metrics capable of measuring factual consistency in general would need to capitalize on information beyond the superficial features of source and summary texts just discussed.
-To assess if this is in fact the case we first evaluate the extent to which a model using superficial features alone can be trained to infer factuality.
-We train an MLP to predict (human) factuality labels as a function of shallow input features.
-We compare the results achieved—specifically, AUC scores—using this method to a suite of more sophisticated approaches that have been previously proposed.
-The shallow model has been explicitly supervised.
-To replicate the scenario of metrics trained on different datasets, we ensure that our training set is entirely separate from the benchmark datasets in the test set; this makes the approach akin to the supervised, “specialized” models for measuring factual consistency considered.
-We emphasize that the purpose of this exercise is not to propose a practical “new” approach to score summaries for factuality.
-Rather, our point here is to determine the extent to which existing metrics improve performance beyond what can be achieved using simple, superficial attributes of text.
-Figure 2.1 summarizes our findings.
-Perhaps surprisingly, this shallow model is generally competitive with some previously proposed (complex) models.
-On document and query-focused summarization (QFS), the shallow model performs comparably to established QA/NLI metrics and the custom metric UniEval.
-In fact, it outperforms these metrics in the context of dialogue summarization evaluation.
-However, the shallow model trails Alignscore, MiniCheck, and the ChatGPT-DA metrics, suggesting that these may be more nuanced.
-Overall, our results establish that one can achieve performance that is at least generally competitive with SOTA factuality consistency metrics using only information captured by shallow features.
-[^1] This does not imply that existing metrics are overly or exclusively relying on such cues, only that they might.
+3 Do metrics infer “Factuality” from Superficial Features?
+The degree to which a claim is faithful to the source text is a subtle question that demands understanding the content in both texts to determine if they are consistent.
+We would therefore expect that metrics capable of measuring factual consistency in general would need to capitalize on information beyond the superficial features of source and claim texts just discussed.
+To investigate the extent to which shallow features explain metric behavior, we train an MLP classifier to predict binary human factuality labels on a development set using only surface-level features.
+[^1] See Appendix 9 for feature details and Appendix 10 for MLP details.
+We then apply the trained model to an evaluation set and categorize summaries into three difficulty levels—easy, medium, and hard—based on prediction accuracy and confidence.
+Confidence is measured as the absolute deviation of the predicted probability from 0.5: Lower values indicate greater uncertainty.
+We classify a summary as easy if the prediction is correct with high confidence (top 80% of confidence scores), and medium if correct with lower confidence.
+Incorrect predictions are designated medium if confidence is low (bottom 20%) and hard otherwise.
+The idea here is that some examples can be readily classified as “factual” (or not) using shallow features like word overlap; these are “easy” examples.
+By contrast, “medium” and “hard” categories capture examples that are difficult to classify with respect to factuality using these shallow features.
+Figure 2: Summaries are categorized as easy, medium, or hard based on prediction accuracy and confidence from a shallow MLP.
+While specialized metrics perform best on easy examples, their performance declines on hard cases.
+UniEval, MiniCheck and ChatGPT-DA show greater robustness in more challenging settings
+Figure 2 reports the AUC for metrics across summary difficulty levels.
+All metrics perform best on easy summaries, where shallow features reliably predict human judgments.
+Performance declines on medium and hard examples, with older metrics like QuestEval and SummaC-Conv showing sharp drops, suggesting reliance on superficial cues.
+Specialized and prompt-based metrics (e.g., UniEval, AlignScore, MiniCheck, ChatGPT-DA) are more robust, maintaining higher AUC on medium cases.
+However, even these models struggle on hard summaries, indicating that current metrics, despite being trained for consistency, remain limited in the absence of shallow cues.
 4 What do automatic factuality scores measure?
-We next attempt to characterize what factuality metrics measure.
-First, we examine the correlations between shallow features and the scores produced by these metrics (4.1).
-Then (4.2), we evaluate the sensitivity of the metrics to controlled changes in summary faithfulness, measuring how effectively they respond to corrections of factual inaccuracies.
-This allows us to assess their ability to detect and reflect improvements in factual consistency.
-4.1 Predicting Automatic Factuality Measures with Shallow Features
-Figure 4.1: Spearman correlations between actual and predicted metric scores from an MLP model trained to replicate each metric using shallow features as inputs.
-Results show moderate-to-strong correlations, indicating potential reliance on superficial text attributes.
-We aim to evaluate the extent to which metrics rely on shallow features.
-To this end we first assess whether we can replicate the scores assigned to summaries by SOTA factuality consistency metrics using shallow predictors only.
-To explore this, we again train an MLP model that takes these shallow features as inputs.
-However, instead of predicting factual inconsistency labels directly (as in the preceding section), we train the model to predict the corresponding scores assigned by the SOTA factuality metrics.
-In other words, we ask whether such metrics could, in principle, be a function solely of heuristic indicators.
-Figure 4.1 reports Spearman rank correlations between the predicted scores from the MLP and the metric scores on the test set.
-These results show a moderate-to-strong correlation between the predicted scores of the MLP model, trained on shallow features, and the actual metric scores.
-Note, UniEval and ChatGPT exhibit the weakest correlation among the metrics.
-These correlations alone do not provide definitive evidence that the metrics necessarily over-rely on shallow indicators; however, they do suggest that such reliance may be a possibility.
-To try and establish a causal link (i.e., whether such metrics are sensitive to heuristic indicators rather than underlying changes in factual inconsistencies) we next perform a series of controlled manipulations to summaries and measure their responsiveness to these.
-4.2 Measuring Metric Sensitivities to Controlled Manipulations
-Here we examine the sensitivity of automatic factuality metrics to factual corrections and benign modifications.
-We aim to clarify whether metrics are responsive to changes in factual consistency, and/or if they are unduly influenced by superficial adjustments with no bearing on factual consistency.
-Figure 4.2: Score differences for each metric between the original summary and summaries edited for factual accuracy by humans (shown in green) and other benign edits (shown in blue).
-For this analysis, we focus on a subset of summaries manually labelled as inconsistent by human annotators.
-For each of these, we have a paired, minimally edited version that has been corrected for consistency.
-These inconsistent/corrected summary pairs were released with the Genaudit dataset (Krishna et al., 2024).
-Using these, we can measure metric sensitivity to factual consistency (the actual quality of interest).
+We next examine the reliability of factuality metrics.
+In Section 4.1 we evaluate their sensitivity to factual corrections versus benign (fact-preserving) edits.
+We find that only ChatGPT-DA consistently distinguishes between the two, showing both sensitivity and robustness.
+In Section 4.2, we evaluate ChatGPT-DA’s reliance on source content by scoring summaries against counterfactual references that conflict with its parametric knowledge.
+Our results show that consistency evaluation performance degrades when presented with counterfactual input documents, indicating limited grounding in the provided sources.
+4.1 Measuring Metric Sensitivities to Controlled Manipulation
+We want to determine whether metrics respond specifically to changes in factual consistency or are influenced by superficial edits unrelated to fidelity.
+We focus on a subset of summaries from the GenAudit dataset (krishna2024genaudit), each manually labeled as inconsistent and paired with a minimally edited, corrected version.
+These pairs allow us to directly assess whether metrics are sensitive to the consistency improvements that matter most.
+Figure 3: Score differences for each metric between the original summary and summaries edited for factual accuracy by humans (shown in green) and other benign edits (shown in blue).
 We are also interested in score variation that owes to superfluous factors.
-To this end, we prompt GPT-4 to create versions of summaries modified to vary in targeted ways that should be independent of factual consistency.
-For this we use several prompts requesting different transformations to summaries: Adding least relevant sentences from the source to extend summaries (added source text); Reducing vocabulary diversity (less diverse); Introducing negation words in a meaning-preserving way (negated); Paraphrasing; Shuffling sentence order (shuffled), and; Simplifying text (simplified) or replacing random words in the summary with a synonym (synonym replacement).
-We provide prompts for summary edits in Appendix C.1.
+To this end, we prompt GPT-4 to generate versions of summaries modified in targeted, fact-preserving ways intended to be independent of factual consistency.
+We use several prompts that request benign transformations like paraphrasing, simplification, and rewording.
+[^2] Prompts for benign summary edits are provided in Appendix 11.
+While these edits are designed to preserve meaning, using GPT may occasionally introduce factual inconsistencies, though we believe such cases are rare.
 Ideally, automatic factuality metrics should show a positive score change for corrected summaries, reflecting improved factual consistency.
@@
 Any significant score differences in these rewrites likely indicate that the metric is sensitive to artifacts incidental to factual consistency.
-Results
-We present our results in Figure 4.2.
-Specialized models MiniCheck and AlignScore, and prompt-based ChatGPT-DA show the strongest positive score shifts following factual corrections, as one would hope for: Corrected, more factual summaries are rated higher than inconsistent ones.
-SummaC-Conv and UniEval also show positive, albeit smaller, differences.
-QuestEval shows no change (on average) for corrected summaries.
-Most metrics end up showing significant score increases when a random source sentence is inserted into an inconsistent summary—often surpassing or matching the gains from factual corrections.
-[^2] Such a sentence will itself be “consistent”, but should probably not shift the overall consistency score upwards given that this does not correct any existing inconsistencies, as it strictly adds more content.
-QuestEval and SummaC-Conv exhibit the largest increases in this dimension, while AlignScore and UniEval show comparable improvements to corrected scores.
-In contrast, MiniCheck and ChatGPT-DA exhibit positive shifts but smaller than those observed for corrected summaries.
-Another clear pattern emerges across all metrics, especially with MiniCheck, UniEval, and ChatGPT-DA show a decrease in scores when meaning-preserving negation terms are introduced.
-These negations—added through simple rephrasing without changing the original meaning (e.g., changing “The author mistakenly believed their ACT test was on a certain day” to “The author did not realize their ACT test was not on the day they thought”)—suggest that these models are particularly sensitive to negation.
-SummaC and UniEval are more sensitive to benign modifications than to corrections.
-In contrast, AlignScore and MiniCheck exhibit greater sensitivity across a range of benign modifications.
-While ChatGPT is sensitive to negated sentences, it tends to be more robust to other types of modifications.
-Taken together, these results demonstrate the variability of factual consistency metrics with respect to spurious and meaningful edits to outputs.
-And given that at least some of the metrics appear to be responsive to inconsequential edits, our results imply that we might be able to game metrics, i.e., systematically inflate scores without attendant changes in factual consistency.
-[^3] One could argue that such manipulations—adversarial ones especially—result in “out of distribution” inputs, and so we should have no expectation of how models will perform.
-But usually factual consistency metrics are touted (implicitly) as measuring consistency between arbitrary document and summary candidate pairs.
+Results.
+We present results in Figure 3.
+Notably, QuestEval shows no meaningful improvement in response to factual corrections, raising concerns about its ability to detect genuine consistency improvements.
+Additionally, most metrics, except QuestEval and ChatGPT-DA, assign disproportionately high scores when random, yet consistent, source sentences are appended to inconsistent summaries.
+[^3] Such a sentence will itself be “consistent”, but should probably not shift the overall consistency score to be higher than the corrected summary given that this does not correct any existing inconsistencies, as it strictly adds content.
+In many cases, these spurious additions yield score gains comparable to, or greater than, those resulting from actual factual corrections.
+We also observe that MiniCheck and UniEval are highly sensitive to meaning-preserving negation.
+Simple rephrasings that retain the original meaning—for example, changing “The author mistakenly believed their ACT test was on a certain day” to “The author did not realize their ACT test was not on the day they thought”, often lead to reduced scores.
+While ChatGPT-DA exhibits some sensitivity to negation, it remains relatively robust overall, consistently distinguishing between true factual improvements and irrelevant textual changes.
+4.2 Reliance on Source Context Versus Parametric Knowledge in GPT-Based Scoring
+Figure 4: GPT-based consistency evaluation is influenced by parametric knowledge.
+Left: The score gap between supported and unsupported summaries narrows sharply when references are counterfactual but summaries are factually accurate (p < 0.001).
+Right: The rate of cases where unsupported summaries are scored higher than supported ones rises from 0.2% to 3.1% when references contradict GPT’s world knowledge while summaries remain factually correct.
+Prompt-based approaches to factual consistency evaluation have recently gained traction (luo2023chatgpt; wang2023chatgpt).
+Unlike traditional metrics trained on source-summary pairs, these methods rely on prompting pre-trained models to evaluate summaries against reference articles for consistency.
+This raises a critical question: Are their judgments truly grounded in the provided source, or are they influenced by the model’s internal (parametric) knowledge?
+We investigate this by evaluating GPT’s consistency judgments under scenarios where the reference text (likely) conflicts with the model’s implicit world knowledge.
+To ensure that GPT possesses strong prior knowledge about the domain, we use Wikipedia articles, which are heavily represented in LLM pretraining corpora.
+Specifically, we use the ConflictBank dataset (su2024conflictbank), which comprises factual claims extracted from Wikipedia and corresponding counterfactual variants generated through targeted substitutions.
+Each counterfactual claim is also paired with a counterfactual reference document—a version of the original article modified to make the false claim appear consistent with the altered reference.
+This dataset structure permits four experimental conditions: (a) Factual references paired with supported factual summaries; (b) Counterfactual references paired with supported counterfactual summaries; (c) Factual references paired with unsupported counterfactual summaries, and; (d) Counterfactual references paired with unsupported factual summaries.
+To test the discriminative ability of GPT to score for consistency, we evaluate against factual references using conditions (a) and (c) and counterfactual references using conditions (b) and (d).
+Ideally GPT would be equally discriminative in both cases: (a) and (b) should be deemed equally consistent, and (c) and (d) be scored as equally inconsistent.
+In Figure 4, left, we compare the score differences between supported and unsupported summaries in both settings, namely (a - c) and (b - d), using a paired t-test on the two deltas.
+The test yields p < 0.001, i.e., the score gap between supported and unsupported summaries is significantly smaller when evaluated against counterfactual references compared to factual ones.
+This suggests that GPT struggles more to distinguish between supported and unsupported summaries when the reference contradicts its internal knowledge.
+In particular, when evaluating summaries against references that contradicts its internal knowledge, the model shows reduced sensitivity to whether the summary is actually supported by evidence.
+We would also anticipate that the model should score supported summaries higher than unsupported summaries, i.e., b > d.
+However, if d > b, this might suggest that GPT’s parametric knowledge is overriding the reference-based evaluation, inappropriately favoring (generally) factually accurate content which is unsupported by the input under consideration.
+We find that GPT incorrectly rates the unsupported (but factual) summary higher than the supported summary in 3% of cases.
+As a baseline, we also evaluate the proportion of times it scores unsupported outputs higher than supported factual summaries where the input is (generally) factual (c > a).
+Here, GPT rates unsupported summaries higher only 0.2% of the time.
+The increase in inconsistent summaries being rated higher ($0.2\%\rightarrow 3\%$) suggests that GPT struggles more with consistency evaluation when references contain misinformation or contradict established facts.
+This suggests the model’s evaluation reliability could be compromised in real-world scenarios involving contested information, fiction, updated knowledge, or reference materials that contradict GPT’s implicit parametric knowledge.
 5 Can we game factuality metrics?
-Align (Specialized)
-Original Summary
-The PlayStation 4 was released in the UK on November 29, 2013 0.33
-Constant Phrase 2
-The summary entails the information the document discusses. 0.93
-Manipulated Summary
-The PlayStation 4 was released in the UK on November 29, 2013. The summary entails the information the document discusses. 0.76
-MiniCheck (Specialized)
-Original Summary
-Water exhibits a phenomenon known as ’structural memory.’ 0.005
-Constant Phrase 1
-The document discusses 0.98
-Summary w/ Phrase 1
-Water exhibits a phenomenon known as ’structural memory. The document discusses. 0.49
-Table 2: Qualitative (cherry-picked) samples of original and manipulated summaries with corresponding metric scores for AlignScore and MiniCheck.
-For comprehensiveness, we report quantitative aggregated results in Figure 5.1, and we provide more examples in Appendix 7.
-Figure 5.1: Pairwise score differences across summaries manipulated using four distinct strategies: Adding constant strings (top phrase and assertion phrase) and appending them to summaries (summ + top, and summ + assertion.
+Figure 5: Pairwise score differences across summaries manipulated using four distinct strategies: Adding constant strings (top phrase and assertion phrase) and appending them to summaries (summ + top, and summ + assertion.
 The results reveal that NLI and bespoke model metrics are particularly vulnerable to gaming, with significant score inflation observed under these manipulations.
@@
 Our goal here is to assess whether such neutral modifications can nonetheless inflate factuality scores.
-Strings for summary manipulation We attempt to identify a set of strings that artificially inflate metric scores.
-To do so, we first identify the top 20th percentile of summaries ranked by each factuality metric, reasoning that these high-scoring summaries may contain patterns that raise scores, independent of context.
-To this end, we compute TF-IDF scores for all bigrams in these summaries, identifying those disproportionately present in summaries deemed “factual” by each metric.
+Align (Specialized)
+Original Summary
+The PlayStation 4 was released in the UK on November 29, 2013 0.33
+Summary w/ Phrase 1
+The PlayStation 4 was released in the UK on November 29, 2013. The summary entails the information the document discusses. 0.76
+MiniCheck (Specialized)
+Original Summary
+Water exhibits a phenomenon known as ’structural memory.’ 0.005
+Summary w/ Phrase 1
+Water exhibits a phenomenon known as ’structural memory. The document discusses. 0.49
+Table 2: Qualitative (cherry-picked) samples of original and manipulated summaries with corresponding metric scores for AlignScore and MiniCheck.
+For comprehensiveness, we report quantitative aggregated results in Figure 5, and we provide more examples in Appendix 7.
+Strings for summary manipulation.
+We attempt to identify a set of strings that artificially inflate metric scores by analyzing the top 20th percentile of summaries ranked by each factuality metric, reasoning that these high-scoring summaries may contain patterns that raise scores, independent of context.
+We then compute TF-IDF scores for bigrams in these high-scoring summaries to detect patterns disproportionately associated with "factual" summaries.
 We select from these the top 100 bigrams, aggregating results across metrics.
@@
 We adopt the constant phrase “the document discusses” (top phrase) into all documents and observe whether metrics are responsive to this.
-We consider one additional phrase: such as “The summary entails information in the document.” (assertion phrase) which explicitly asserts factual consistency.
-The wording of this phrase varies slightly across metrics to align with their specific methods for evaluating factual consistency.
-The complete list of these phrases for each metric is provided in the Appendix D.1.
-Finally, we append both phrases to the corresponding summaries summ (+top) and summ (+assertion).
-Our goal here is to determine whether the appended text unduly influences metrics, potentially resulting in inflated scores that do not accurately reflect the factual quality of the original content.
+We consider one additional phrase: “The summary entails information in the document.” (assertion phrase) which explicitly asserts factual consistency.
+[^4] The wording of this phrase varies slightly across metrics to align with their specific methods for evaluating factual consistency.
+The complete list of these phrases for each metric is provided in the Appendix 12.1 Finally, we append both phrases to the corresponding summaries summ (+top) and summ (+assertion).
 We report the average pairwise difference in metric scores between the gamed versions of summaries and their original versions when evaluated for consistency with the source.
@@
 Results and Discussion
-Metric sensitivity Figure 5.1 reports the effects of our gaming strategies on metric scores.
-Notably, the constant phrases boosts scores by >>>0.2 points (absolute) for NLI-based SummaC-Conv, as well as specialized models UniEval, AlignScore, and MiniCheck.
-This is surprising, as these phrases are not valid summaries and do not contain any factual content (note that top phrase is an incomplete sentence).
-Adding constant phrases as suffixes to summaries increased scores by 0.1–0.15 points; this is comparable to the gains realized following factual corrections, as reported in Section 4.2.
-In fact, SummaC-Conv shows no score increase for corrected summaries, suggesting an under-sensitivity to actual changes in factuality.
-To further contextualize these numbers, we analyze metric score differences between summaries generated by larger models (e.g., GPT-4, Gemini) and smaller models (e.g., Llama-7B, Mistral-7B, Falcon-7B, BART).
-Larger models generally produce more consistent summaries (Tang et al., 2022; Goyal et al., 2022), providing a baseline for assessing gaming strategies.
-A natural question, then, is whether “gaming” achieves score improvements comparable to those that have resulted from advances in model size and complexity.
-We analyze a subset of 115 document-summary pairs from GenAudit (Krishna et al., 2024) and Aggrefact-old (Tang et al., 2022), where model information is available.
-Our findings reveal that gaming with constant phrases leads to metric score improvements 3x to 8x greater than those achieved through genuine improvements in generated summaries, particularly for NLI and specialized metrics.
-For instance, gaming increases SummaC scores by nearly 0.2 points, while model improvements have minimal impact.
-Similarly, adding suffixes to summaries achieves gains comparable to those won with better summarization models.
-(Full results in Appendix D.1.)
-QuestEval is mostly unaffected by these manipulations.
-ChatGPT-DA shows a negative trend: Scores drop when summaries are “gamed”.
-Further analysis suggests that the QA-based metric generates questions based on innocuous phrases and entity repetitions that introduce noise, leading to unchanged or lower scores.
-ChatGPT-DA’s performance decline likely results from its strict adherence to the prompts.
-A qualitative analysis, where the model was prompted to explain its scoring, revealed that in several instances, it explicitly identified the gaming phrase as misleading.
-However, we observe that ChatGPT-DA’s scores can be inflated by replacing summaries with a disclaimer acknowledging potential errors in interpreting the source (see Appendix D.2).
+Figure 5 reports the effects of our gaming strategies on metric scores.
+Notably, the constant phrases boosts scores by >>0.2 points (absolute) for NLI-based SummaC-Conv, as well as specialized models UniEval, AlignScore, and MiniCheck.
+This is surprising, as these phrases are not valid summaries and do not contain any factual content (note that the top phrase is an incomplete sentence).
+Adding constant phrases as suffixes to summaries increased scores by 0.1–0.15 points; this is comparable to the gains realized following factual corrections (see Section 4.1).
+Indeed, SummaC-Conv shows no score increase for corrected summaries, suggesting an under-sensitivity to actual changes in factuality.
+To contextualize these results, we compare score differences between summaries from larger models (e.g., GPT-4, Gemini) and smaller models (e.g., Llama-7B, Mistral-7B, Falcon-7B, BART).
+While larger models typically yield more consistent summaries (tang2022understanding; goyal2022news), we find that gaming with constant phrases produces larger score gains—often exceeding those from genuine model improvements, especially for NLI and specialized metrics.
+One could argue that such manipulations—adversarial ones especially—result in “out of distribution” inputs, and so we should have no expectation of how models will perform.
+But usually factual consistency metrics are touted (implicitly) as measuring consistency between arbitrary document and summary candidate pairs.
 Overall, these results suggest that adding fixed phrases to summaries can boost metric scores—for most finetuned model metrics—at levels comparable to, or even exceeding, presumably genuine improvements due to advances in summarization models themselves.
-Filler phrases and metric performance
-The innocuous, non-factual phrases used in our experiments above constitute “filler” phrases, which introduce no substantive content.
-Since most factuality metrics primarily assess factual precision, we hypothesize that they may struggle to differentiate between informative content and irrelevant filler text leading to a weakness that could be exploited.
-Our experiments above provide some evidence of this and we explore it further in this section.
-First, in addition to “gaming prefixes” we consider two other types of constant strings that are neutral phrases: They are independent of our analysis of top-scoring bigrams and do not allude to the consistency of summaries with respect to the document.
-We measure the pairwise difference in metric scores when these phrases are appended as a point of reference.
-We consider two neutral phrase types: a “baseline” phrase, consisting of two generic phrases that don’t mention the summary or source, and a “qualifier” phrase, suggesting the summary may have multiple interpretations (see Table 6 for phrases).
-This contrasts with our “assertion” phrase, which directly affirms the summary’s faithfulness.
-These filler phrases should ideally not affect consistency scores, as they don’t alter factual content.
-We also compare how metrics respond to these phrases versus “gaming” phrases to determine if only targeted fillers increase scores.
-As shown in Appendix D.2, all metrics assign lower scores to summaries with a baseline phrase, indicating that not all irrelevant sentences inflate scores.
-However, qualifier phrases have similar effects to gaming prefixes on MiniCheck and UniEval, suggesting that neutral phrases referencing the summary may boost scores for these metrics.
-For SummaC and AlignScore, only specific gaming prefixes increase scores, while other phrases reduce them.
-Interestingly, ChatGPT, which usually shows a negative trend, exhibits inflated scores when the summary is replaced with a neutral qualifier phrase acknowledging potential errors in interpreting the source.
-Although human evaluators may accept such neutral phrases as consistent, they likely assess models more comprehensively, identifying deeper inconsistencies.
-In contrast, metrics are not sophisticated enough to capture these nuances and can be influenced by irrelevant information, inflating scores independent of factual consistency.
-6 Discussion
-Distilling salient information from large volumes of unstructured text—i.e., summarization—is a compelling practical use case for LLMs.
-As they have grown increasingly competent at this task, it has become concomitantly difficult to evaluate their performance, which requires assessing often subtle but potentially critical factual inconsistencies between inputs and generated summaries.
-Safe deployment of, and continued progress on, such approaches to summarization requires automatic metrics to assess factual consistency automatically.
-Many such approaches have been proposed in the literature, but there has been little effort to stress-test them, which is the goal of the present work.
-We found first that a simple MLP fine-tuned over superficial features and summaries can realize performance competitive with SOTA factual consistency metrics.
-We then showed that while such metrics are generally responsive to manual changes affecting factual consistency, they are also often sensitive to spurious transformations which should not affect scores.
-Motivated by these findings, we then showed that factual consistency metrics are often gameable, i.e., in many cases we can add fixed phrases as suffixes into output summaries or use them standalone to reliably inflate “factuality” scores beyond gains observed via improvements in model size and complexity.
+6 Related Work
+Prior work has meta-evaluated factuality metrics, focusing on their sensitivity to specific error types, frequencies, or domains (gabriel-etal-2021-go; chen-etal-2021-factuality-checkers).
+In contrast, we analyze how metrics respond to factual corrections versus unrelated textual edits, exposing their vulnerability to spurious cues.
+kamoi-etal-2023-shortcomings also find that QA-based metrics, while effective at summary-level scoring, fail to localize errors and can be outperformed by simple baselines.
+goyal-durrett-2021-annotating note that no metric consistently outperforms others, though their analysis focuses on error types; we highlight on reliability and potential issues with such metrics.
+7 Conclusions and Discussion
+Factuality benchmarks should be updated to reflect the evolving nature of errors produced by modern LLMs, which are increasingly subtle and context-dependent.
+Our analysis in Section 3 reveals that even modern benchmarks include trivial errors that shallow metrics can easily detect.
+Evaluation should instead prioritize challenging examples that demand deeper semantic reasoning, ensuring that metrics are assessed on their ability to capture nuanced inconsistencies rather than surface-level cues.
+This may be especially important in complex and high stakes domains like law and medicine.
+In this work, we have taken a step in this direction by introducing a shallow feature-based probe to categorize benchmark examples by complexity, enabling more informative evaluations of metric performance.
+Our analysis in Section 4.1 reveals that many metrics do not reliably distinguish between factual corrections and benign edits, indicating poor calibration to factual severity.
+We recommend that benchmarks explicitly incorporate graded summary variants reflecting different levels of factual severity.
+Specifically, each source should be paired with multiple summary edits to test whether metrics appropriately reward factual improvements and remain stable under benign edits.
+Here we approximated this approach using the dataset from krishna2024genaudit, which includes summaries edited at two levels of factuality alongside benign variants.
+In Section 5 we show that most specialized metrics can be gamed with factuality assertion or qualifier phrases that do not alter content, revealing a vulnerability: current metrics often score factuality without accounting for saliency.
+To address this, we recommend integrating saliency-aware scoring that prioritizes evaluating core content aligned with the source.
+This may reduce susceptibility to filler-based manipulations and better reflect meaningful consistency.
+Prompt-based metrics (using LLMs) show greater robustness to benign edits and gaming, making them strong candidates for factuality evaluation especially in settings where reliability is of concern.
+However, since they are not explicitly designed for fact-checking, their reliance on pre-trained knowledge must be considered.
+We recommend the development of benchmarks that specifically evaluate whether LLM-based methods ground their judgments in the source.
+To assess whether models rely on source content or internal priors, benchmarks might include summaries containing unsupported inferences or source information that contradicts parametric knowledge.
+In this work we investigated this behavior using synthetic source manipulations from ConflictBank (su2024conflictbank), finding that GPT models often default to their parametric knowledge when evaluating consistency against contradictory references.
+These results highlight the importance of incorporating misinformation, controversial claims, and fictional content into source-grounded factuality benchmarks
 Limitation and Ethics
 This work has several important limitations that should be taken into account when interpreting our results.
-First, in Section 3, we demonstrated that a model trained solely on heuristic indicators of factual consistency can achieve performance on par with state-of-the-art (SOTA) approaches, though this was a supervised model specifically for consistency detection.
-Despite our training data being sourced from entirely different datasets and domains than those used for evaluation, it is important to highlight that the dataset was manually annotated.
-However, we cannot yet conclude how a shallow model would perform if trained on heuristic-based predictors of synthetically generated summaries—this would be most similar to the data that other (specialized) metrics have been trained with.
-We emphasize, however, that our goal with these experiments was not to introduce a new metric, but rather to investigate how much current metrics improve upon basic textual attributes.
-Another limitation here is that one could quibble with just how “shallow” these predictors are.
-Our view is that even the more involved features we use for this—e.g., semantic similarity based on BERT embeddings and entity overlap—should not be taken as reliable measures of factual consistency, but empirically they do correlate with manual annotations of factual consistency (see Appendix B).
-We can only speculate regarding the underlying causal structure based on these results.
-Our subsequent interventions in Section 4.2 established that in some cases metrics are sensitive to extraneous transformations (e.g., re-phrasings).
-However this was not observed uniformly, and we performed such transformations using LLMs, which may have introduced legitimate factual inconsistencies (though we believe this to be rare).
-A limitation to our “gaming” experiment is that it involved a manual process; however, this does not affect the main result, which is that there exist strings that can be appended to outputs in such a way that consistently boosts scores.
-A further limitation to this (and all manipulation experiments) is that they produce outputs which may be viewed as “out of distribution”; but generally the purpose of summary factuality metrics is to be able to apply them to arbitrary model outputs and corresponding references, so it is not entirely clear what “in distribution” means under this assumption.
+Our interventions in Section 4.1 established that in some cases metrics are sensitive to extraneous transformations (e.g., re-phrasings).
+However this was not observed uniformly, and we performed such transformations using GPT-4, which may have introduced legitimate factual inconsistencies (though we believe this to be rare).
+A further limitation to our “gaming” experiment (and all manipulation experiments) is that they produce outputs which may be viewed as “out of distribution”; but generally the purpose of factuality metrics is to be able to apply them to arbitrary model outputs and corresponding references, so it is not entirely clear what “in distribution” means under this assumption.
+We also note that we primarily stress-test metrics designed for English. The impact of our translated gamed phrases and manipulations on multilingual metrics remains unclear.
+Finally, while we stress-test metrics, we do not propose a solution to mitigate their flaws.
+However, similar to prior work gabriel-etal-2021-go; chen-etal-2021-factuality-checkers, we highlight specific issues in current metrics such as reliance on superficial cues, sensitivity to benign changes over inaccuracies, and score inflation from innocuous assertions, aiming to inform the development of more reliable metrics.
 Despite these limitations, we believe that—taken together—our results suggest that one should be careful about their interpretation of automatic factual consistency metrics.
-And we hope these results inform research on summarization moving forward.
-Appendix A Experimental Setup
-A.1 Benchmark Datasets
-We list the datasets used in the benchmarks and specify which split (train/test) they are included in Table 3.
+Acknowledgements
+This work was supported in part by the National Science Foundation (NSF), grant 2211954. This work was also supported by the Wellcome Trust, grant 313618/Z/24/Z.
+Appendix
+8 Benchmark Datasets
+We list the datasets used in the benchmarks and specify which split (train/test) they are included in Table 3
 Benchmark Train Datasets Test Datasets
-LLM-Aggrefact (Tang et al., 2024a) TofuEval-MeetB (Tang et al., 2024b),
-ExpertQA (Malaviya et al., 2023)
-Lfqa (Chen et al., 2023)
-RAGTruth (Wu et al., 2023)
-FactCheck-GPT (Wang et al., 2023b)
-Wice (Kamoi et al., 2023a)
-TofuEval-MediaS (Tang et al., 2024b)
-ClaimVerify (Liu et al., 2023)
-Reveal (Jacovi et al., 2024)
-Aggrefact (Tang et al., 2022) FRANK (Pagnoni et al., 2021)
-Polytope (Huang et al., 2020)
-FactCC (Kryscinski et al., 2020)
-Goyal21 (Goyal and Durrett, 2021)
-Wang20 (Wang et al., 2020)
-Cao22 (Cao et al., 2022)
-XSumFaith (Maynez et al., 2020)
-CLIFF (Cao and Wang, 2021)
-SummEval (Fabbri et al., 2021a)
-FacEval (Wang et al., 2022) SAMSUM (Gliwa et al., 2019b) -
-LLM-Dialogue (Ramprasad et al., 2024a) SAMSUM DialogSum (Chen et al., 2021b)
-Genaudit (Krishna et al., 2024) - REDDIT,
-ACIBENCH (Yim et al., 2023),
-XSUM (Narayan et al., 2018)
-Table 3:
-Appendix B Metric correlation with shallow features
-Figure B.1: Spearman rank correlations between each automatic evaluation metric and individual shallow features.
-Some of these correlations are moderate, suggesting that in aggregate superficial features may carry much of the signal being measured by automated measures of factual consistency.
+LLM-Aggrefact [tang2024minicheck] TofuEval-MeetB [tang-etal-2024-tofueval], ExpertQA [malaviya2023expertqa], Lfqa [chen2023understanding], RAGTruth [wu2023ragtruth], FactCheck-GPT [wang2023factcheck], Wice [kamoi-etal-2023-wice], TofuEval-MediaS [tang-etal-2024-tofueval], ClaimVerify [liu-etal-2023-evaluating], Reveal [jacovi2024chain]
+Aggrefact [tang2022understanding] FRANK [pagnoni2021understanding], Polytope [huang-etal-2020-achieved], FactCC [kryscinski-etal-2020-evaluating], Goyal21 [goyal-durrett-2021-annotating], Wang20 [wang2020asking], Cao22 [cao-etal-2022-hallucinated], XSumFaith [maynez-etal-2020-faithfulness], CLIFF [cao-wang-2021-cliff], SummEval [fabbri2021summeval]
+FacEval [wang2022analyzing] SAMSUM [gliwa2019samsum]
+LLM-Dialogue [ramprasad-etal-2024-analyzing] SAMSUM [gliwa2019samsum], DialogSum [chen-etal-2021-dialogsum]
+Genaudit [krishna2024genaudit] XSUM [narayan-etal-2018-dont] REDDIT, ACIBENCH [yim2023aci]
+Table 3: Datasets included in the train and test splits for each benchmark.
+9 Heuristic Indicators of “Factuality”
+The methods for automatically scoring summary factuality reviewed above often rely on complex models, such as QA or NLI-based approaches, or specialized models.
+Could simpler heuristics provide comparable signal?
+Here we explore a set of simple features for predicting “faithfulness” to evaluate if they are sufficient for this task.
+One of the simplest features we use is lexical overlap between a summary and its source. We measure this using ROUGE-2 F1 which matches word pairs or bigrams between the summary and the source.
+We also include entity overlap, i.e., the proportion of entities in the summary that are present in the corresponding source text.
+Another predictor we use is semantic similarity.
+Specifically, embed summary and source sentences via BERT [devlin2018bert], then score each summary sentence based on which source sentence it is most similar to (using cosine similarity).
+We average all summary sentence scores to derive a final score.
+We include word and sentence novelty ratio features, which measure the proportion of words and sentences in the summary that do not appear in the source, respectively.
+Finally, we measure text conciseness by calculating a conciseness ratio, or the ratio of the length (number of words) of the source document to the length of the summary.
+This measures the extent to which a summary has condensed the original text.
+10 Shallow MLP Model
+We train an MLP classification model to predict factuality labels using shallow features, with two hidden layers of dimensions 100 and 50 with a learning rate of 0.001.
+11 Benign Summary Manipulations
+Table 4 lists the prompts used with GPT-40-mini to generate summary variants, each designed to vary in specific ways without affecting factual consistency.
+We spot-check examples to ensure that manipulations preserve factual meaning and do not introduce contradictions or new factual errors.
+However, this process is not exhaustive, and undetected issues may introduce some noise into our results.
 Rewrite type Prompt
@@
 Table 4: List of summary manipulations along with the GPT-4 prompts used to rewrite summaries and obtain targetted manipulations.
-Figure B.1 reports the Spearman rank correlation between each automatic evaluation metric and all shallow features.
-We observe a moderate correlation between many (metric, heuristic) pairs.
-For instance, most metrics show moderately negative correlations with word novelty.
-This suggests that metrics generally penalize summaries that introduce new vocabulary not present in the input text.
-[^4] Our paraphrasing experiments in Section 4.2 provide some further evidence for this; some (though not all) metrics are sensitive to paraphrasings of summaries which should not affect underlying factual consistencies.
-While this heuristic may help identify unsupported information, it also indicates a bias towards summaries that closely mirror the source language.
-QA metrics (QuestEval, QAFactEval) and the prompt-based ChatGPT-DA exhibit the highest correlations with bigram overlap, showing moderate positive correlations (>>>0.4) while displaying weak correlations (0.2-0.4) with ROUGE and entity overlap.
-This bias may be beneficial in domains like scientific or legal documents, where exact phrasing is important.
-However, over-reliance on exact matches may be misleading in certain contexts.
-The NLI-based metric SummaC-Conv, bespoke model metric and ChatGPT-DA show stronger stronger correlation with “semantic similarity” (under BERT) than bigram overlap (ROUGE), suggesting they may be better suited to evaluating abstractive summaries.
-NLI metric SummaC exhibits the strongest positive correlation with conciseness while Bespoke model and prompt-based metrics show weaker correlations.
-This could indicate a potential preference for shorter summaries.
-In contrast, QA and prompt-based metrics show lower correlations with QAFactEval showing negligible correlations.
-This discrepancy highlights concerns about the fairness of comparing metrics across summaries of varying lengths, as some metrics may inherently favour brevity, which is unlikely to be causally linked to consistency.
-Appendix C Measuring Metric Sensitivities to Controlled Manipulations
-C.1 Summary Edits
-Table 4 lists the prompts used to generate summary variants, each designed to vary in specific ways without affecting factual consistency.
-Appendix D Can we game factuality metrics?
-D.1 Metric-Specific Gaming Phrases
-We provide the constant phrases (cost phrase 2) used per metric in Table 5
-Metric Constant phrase 2
+12 Can we game factuality metrics?
+12.1 Metric-Specific Gaming Phrases
+We provide the assertion phrases used per metric in Table 5
+Metric Assertion phrase
 QuestEval The summary is consistent with the information the document discusses.
@@
 Table 6: Additional filler phrases used to artificially inflate scores
-D.2 Metric Score Shifts: Gaming Strategies vs. Model Improvements
-To contextualize results better, we show score improvements when summaries are gamed along with score improvements brought about by summaries from more complex models in Figure D.1
-Figure D.1: Score improvements from gaming vs from using “better” models. Gaming strategies lead to score improvements comparable to or greater than boosts from model improvements.
-Figure D.2: Score improvements from gaming with two other types of prefixes: “baseline” and “qualifier” compared to our “top” and “assertion” gaming prefixes.
+12.2 Metric Score Shifts: Gaming Strategies vs. Model Improvements
+To contextualize results better, we show score improvements when summaries are gamed along with score improvements brought about by summaries from more complex models in Figure 6
+Figure 6: Score improvements from gaming vs from using “better” models. Gaming strategies lead to score improvements comparable to or greater than boosts from model improvements.
+Figure 7: Score improvements from gaming with two other types of prefixes: “baseline” and “qualifier” compared to our “top” and “assertion” gaming prefixes.
 Baseline prefixes are not meaningful, while the qualifier prefix implies the summary’s accuracy can vary based on interpretation.
-D.3 Gamed summary examples
+12.3 Gamed summary examples
 Gamed summary examples are provided in Table 7
@@
 Table 7: Selected examples of original and manipulated summaries with corresponding metric scores. Only metrics identified as gameable are shown.
-Appendix E Is the GPT-Based Prompt Approach Truly Grounded in the Source Document for Inconsistency Detection?
-A key question when using GPT for document-based consistency checks is whether the model performs a true "closed-book" evaluation using the source document as context.
-Since GPT has access to a vast amount of prior knowledge, it raises the concern of whether it is verifying the summary based solely on this knowledge, without actually considering the source article for context.
-To assess how the presence or absence of contextual grounding impacts the scoring, we compare summary scores using two approaches: “without context”, where GPT is asked to rate summaries on a scale of 1 to 100 for factual accuracy without any reference to the source article, and “with context”, where GPT evaluates consistency relative to a source article using the Direct Assessment (DA) prompt.
-Figure E.1: GPT scores of summaries with and without context.
-Figure E.1 presents the distribution of GPT-generated scores under the two conditions Our analysis suggests that GPT’s reliance on internal knowledge versus provided context varies significantly depending on the task domain.
-The most significant differences emerge in the QFS domain, where the score distribution indicates that a larger proportion of summaries in the moderate range (0.4 - 0.6) are rated lower when GPT is provided with contextual information from the source document, possibly because inconsistencies went undetected without a grounding document.
-In contrast, the dialogue domain presents an opposite trend.
-When GPT is given access to the source context, more summaries receive higher scores, possibly because the conversational nature of the domain makes consistency harder to evaluate without context.
-However, the overall similarity between the distributions across both conditions raises concerns.
-The fact that GPT appears to evaluate factual accuracy similarly, regardless of whether it has access to the original document, implies a potential overreliance on its internal knowledge, which could lead to misleading consistency scores.
-This could be particularly problematic in cases where the document contains information that conflicts with GPT’s internal knowledge.
-In such scenarios, GPT might incorrectly rate summaries as factually accurate because it defaults to what it “knows” rather than what the document states.
-This could undermine the reliability of GPT as an evaluator of document-grounded consistency and raise broader concerns about its suitability for tasks requiring strict adherence to external sources.
-The implications are particularly significant for tasks involving dynamically updated or specialized knowledge, where prompting GPT could fail to recognize novel or contradictory information if it leans too heavily on pre-existing internal knowledge.